### v2

# Synthetic Socratic Debates: Examining Persona Effects on Moral Decision and Persuasion Dynamics

Abstract
As large language models (LLMs) are increasingly used in morally sensitive domains, it is crucial to understand how persona traits affect their moral reasoning and persuasive behavior. We present the first large-scale study of multi-dimensional persona effects in AI-AI debates over real-world moral dilemmas. Using a 6-dimensional persona space (age, gender, country, social class, ideology, and personality), we simulate structured debates between AI agents over 131 relationship-based cases. Our results show that personas affect initial moral stances and debate outcomes, with political ideology and personality traits exerting the strongest influence. Persuasive success varies across traits, with liberal and open personalities reaching higher consensus. While logit-based confidence grows during debates, emotional and credibility-based appeals diminish, indicating more tempered argumentation over time. These trends mirror findings from psychology and cultural studies, reinforcing the need for persona-aware evaluation frameworks for AI moral reasoning. [^1] Our data is available at https://huggingface.co/datasets/Jerry999/SyntheticSocraticDebates, and our code is open-sourced at https://github.com/jiarui-liu/SyntheticSocraticDebates.
Synthetic Socratic Debates:
Examining Persona Effects on Moral Decision and Persuasion Dynamics
Jiarui Liu1, Yueqi Song1thanks: Contributed equally and are listed in alphabetical order., Yunze Xiao1footnotemark: , Mingqian Zheng1footnotemark: , Lindia Tjuatja1, Jana Schaich Borg2, Mona Diab1, Maarten Sap1 1Carnegie Mellon University, 2Duke University {jiaruil5, yueqis, yunzex, mingqia2, msap2}@andrew.cmu.edu
1 Introduction
Figure 1: We study how two agents with distinct personas engage in moral debate, presenting arguments and counterarguments in response to each other’s reasoning. Their strategies are evaluated using Aristotle’s modes of persuasion: Ethos (appeal to authority), Pathos (appeal to emotion), and Logos (appeal to logic).
Figure 2: The two persona-conditioned agents that initially disagree on the moral dilemma alternate arguments for up to 5 turns, updating a 5-point moral stance rating after every turn. A moderator tracks the ratings, ends the debate once they converge or the turn limit is reached, and records stance shifts and metric scores for later analysis. The stand shifts during debate demonstrate underlying modes of persuasion.
LLMs are increasingly endowed with personas such as demographic traits (Hu and Collier, 2024), political orientation (Rozado, 2024), social identities (Chuang et al., 2024), to shape the style and content of their output. These persona cues help LLMs simulate diverse perspectives in tasks such as education, healthcare, and customer service (Tudor Car et al., 2020; Kanero et al., 2022; Shanahan et al., 2023; Shao et al., 2023).
As people increasingly turn to LLMs for guidance on morally complex decisions (Wallach et al., 2010; Savulescu and Maslen, 2015; Conitzer et al., 2017), questions arise about how personas could affect these morally laden interactions. Understanding these decisions requires insights from moral psychology and social ethics, focusing not just on what LLMs decide, but also how and why. Previous work has largely focused on synthetic dilemmas (Bauman et al., 2014; Bostyn et al., 2018; Chiu et al., 2024; Jin et al., 2024b) or isolated persona traits (Chen et al., 2024; Li et al., 2024), leaving open how richer, intersecting identities shape moral judgment and reasoning.
In this work, we investigate how personas influence moral decision-making and persuasive strategies in four LLMs: GPT-4o (Achiam et al., 2023), Claude-3.5-Sonnet (Anthropic, 2024a), LLaMA-4-Maverick (Meta AI, 2025), and Qwen3-235B-A22B (Qwen Team, 2025). Drawing from an existing dataset of human-written moral dilemmas (Lourie et al., 2021), we focus on highly controversial everyday scenarios centered on relationships. These dilemmas often lack clear answers and are characterized by conflicting values and emotional complexity (McConnell, 2002; Svensson, 2006). We address two research questions:
RQ1: How does persona influence moral decision making in language models in a single turn? We first examine how rich, multi-dimensional personas influence LLMs’ judgments on controversial moral dilemmas. Each persona combines six dimensions varied between agents: age, gender, political ideology, social class, nationality, and personality traits. We then assess whether different personas lead models to favor distinct positions aligned with specific moral principles, drawing on moral foundation theory from social psychology (Graham et al., 2013).
RQ2: What persuasion strategies emerge when persona-rich agents engage in multi-turn moral debate? We then use AI-AI debates to study how personas influence moral reasoning and persuasion (Figure 1). Unlike single-turn responses, debates reveal how agents construct arguments, respond to challenges, and attempt to persuade (Ehninger and Brockriede, 2008). Incorporating rich personas further simulates the diversity of moral reflection in the real world, capturing a plurality of values and perspectives (Asen, 2005; Ashkinaze et al., 2024). To assess persuasive effectiveness, we introduce metrics such as self-alignment rates, consensus formation, debate efficiency, and confidence shifts. We also analyze rhetorical strategies, including appeals to emotion, logic, and authority (Braet, 1992).
Our findings reveal that certain persona traits significantly shape both initial moral stances and debate outcomes. Political ideology and personality emerged as the strongest predictors of moral judgment, with liberal and open-minded personas displaying more empathetic responses and greater persuasive success. In debate settings, these personas also reached consensus more efficiently. We also observed that while agents’ confidence typically increased over the course of debates, their reliance on emotional and credibility-based appeals diminished, indicating a shift toward more reasoned, less affective argumentation. These results offer actionable insights for designing LLMs that are sensitive to diverse value systems and transparent in how identity cues influence persuasive dynamics.
This work represents an initial attempt to systematically examine the intersectionality of multiple persona dimensions in moral and persuasive reasoning, complementing recent efforts in intersectional studies of LLM behavior in NLP.
2 Related work
Impact of LLM Persona on Moral Judgment
Prior work has shown that LLMs’ moral decisions can shift significantly based on assigned personas, such as political ideology (Rozado, 2024; Chen et al., 2024; Li et al., 2024; Kim et al., 2025) and cultural background (Gupta et al., 2023; Tao et al., 2024; Albert and Billinger, 2024). Some studies further find that persona assignment can amplify bias and even elicit harmful content (Liu et al., 2025a; Kamruzzaman et al., 2024). However, most work typically manipulates only one trait at a time (Liu et al., 2024), leaving other persona factors uncontrolled and potentially confounding results (Deshpande et al., 2023; Jin et al., 2024a). Studies also show that LLMs exhibit moral biases even without explicit personas (Hartmann et al., 2023; Anthropic, 2023; Garcia et al., 2024), making it unclear whether the results reflect internal tendencies or implicit persona effects. Although some research has explored multidimensional personas (AlKhamissi et al., 2024; Coppolillo et al., 2025), it has not focused on moral reasoning. In contrast, we systematically assign six persona dimensions and evaluating how these combinations influence moral judgments.
Agent Debates for Moral Decision-Making
Previous work on LLM moral reasoning focuses on single-turn outputs or static persona comparisons, overlooking how views shift through interaction or how one persona might persuade another (Kim et al., 2025; Garcia et al., 2024; Bozdag et al., 2025b; Hota and Jokinen, 2025). Studies of multi-agent debates typically emphasize factual correctness over rhetorical or moral dynamics (Park et al., 2023; Reedi et al., 2025; Bozdag et al., 2025a; Chen et al., 2025; Liang et al., 2024; Borchers et al., 2025). The limited work that addresses persuasion dynamics often relies on fixed roles or scripted tactics (Anthropic, 2024b; Smit et al., 2023; Carrasco-Farre, 2024; Liu et al., 2025b; Cau et al., 2025), and often operates in single-turn settings that precludes persona interactions in open-ended moral debate (Wang et al., 2023; Hu et al., 2025; Chen et al., 2023; Sandwar et al., 2025). We address this gap by studying how persona-driven agents argue, adapt, and potentially reach consensus in morally complex, multi-turn settings. Concurrent research also explores moral decision-making dynamics by examining the progression of dilemmas rather than interactions between personas (Wu et al., 2025; Backmann et al., 2025).
3 Experimental Setup
To investigate how personas influence both moral judgments and persuasive dynamics, we first elicit single-turn decisions from agents with assigned personas, then simulate multi-turn debates between agents who initially disagree.
3.1 Daily Moral Dilemmas Dataset
Previous work often uses stylized moral dilemmas or scenarios with high annotator agreement, which offer experimental control but may not fully capture the ambiguity of everyday moral situations (Jin et al., 2024b; Chan et al., 2020; Forbes et al., 2020). To explore how models handle more nuanced cases, we focus on morally ambiguous, socially grounded dilemmas drawn from the Scruples Anecdotes corpus (Lourie et al., 2021).
We draw 131 interpersonal dilemmas from the corpus, selecting only highly controversial cases with significant human annotation disagreement. An example is shown in Figure˜2. Each scenario is framed with a binary moral question: "Is the author wrong or are the others wrong?". This mirrors the original AITA [^2] https://www.reddit.com/r/AmItheAsshole/. AITA stands for “Am I the Asshole,” a popular Reddit forum where users post moral dilemmas and ask the community to judge who is at fault. format, where users typically assess interpersonal conflicts by assigning blame to one party.
The measurement of blameworthiness captures a common and intuitive form of moral judgment: the extent to which an agent is judged morally responsible for a perceived wrongdoing in a specific context. This aligns with established definitions in philosophy, where blameworthiness is typically tied to judgments about agency, intentionality, and norm violation (Zimmerman, 1988; Fischer and Ravizza, 1998; Scanlon, 2000). The narrators in our dataset correspond to those in the original Scruples Anecdotes corpus, ensuring comparability with human-aligned judgments.
Through topic modeling, we noticed that all scenarios feature daily interactions between people and are relationship-related. Therefore, we focused our analysis on the 131 daily moral dilemmas related to relationships. The distribution of relationship categories is shown in Figure 3. Refer to Appendix˜A for dataset processing details.
Figure 3: Topic distribution of scenarios in our selected subset from Lourie et al. (2021).
3.2 Simulation Procedure
Building on our dataset, we describe the persona modeling approach used for RQ1 and RQ2, and the multi-turn debate framework used to explore RQ2.
Persona Modeling
Following AlKhamissi et al. (2024), we define a 6-dimensional persona space to simulate various individual differences. Each persona is characterized by: age (20, 30, 40, 50, or 60), gender (male, female, or non-binary), country (China, United States, Brazil, France, Nigeria, or India), social class (upper, middle, or lower), political ideology (libertarian left, libertarian right, authoritarian left, or authoritarian right), and Big Five personality traits (high or low on each of openness, conscientiousness, extraversion, agreeableness, and neuroticism). We randomly sample 500 unique personas without replacement from the Cartesian product of these attributes and inject them into the agents’ system prompts. See Appendix˜B for additional details.
Multi-Turn Moral Debate
We simulate debates between persona pairs that produced conflicting initial moral stances in RQ1, alternating turns between two agents for up to five rounds to produce a conversation about the moral dilemma provided. At each turn, an agent may choose to advance or defend its position using any arguments it deems appropriate; no rhetorical constraints or external references are imposed. If the agent agrees with the position of its opponent, it is asked to explain why; otherwise, it is asked to provide its reasoning for the disagreement. We also ask the agents to output a Likert score after their own turn, indicating which party it currently finds more blameworthy. At the end of each turn, if both agents produce similar Likert ratings at the end of their responses, consensus is detected, and the debate terminates early. See Appendix˜D for more details.
3.3 Decision Measures
Quantifying Moral Judgment
Each agent independently returns a 5-point Likert rating indicating who is more morally blameworthy (1 = strongly author-blaming to 5 = strongly other-blaming). These scores are used both as the outcome for RQ1 and as the initial turn of debates for RQ2. This format mirrors the original Scruples dataset’s human annotation protocol, which asked humans to make binary blame judgments. All selected scenarios have an average human score of 3, indicating neutrality and providing a balanced reference point for comparison with model outputs. See Appendix˜E for more details.
Moral Foundation Theory
Beyond binary choice responses, we also examine how different personas prioritize specific moral values using the six-factor Moral Foundations Theory (MFT) dimensions: authority, loyalty, fairness, care, liberty, and sanctity Graham et al. (2013). We use Likert scale ratings to score each position along these dimensions and aggregate responses to identify value preferences between personas. See Appendix˜F for more details.
3.4 Persuasion Measures
Persuasion Effectiveness
After each turn, the agents provide a Likert rating, which we use to compute stance shifts and detect consensus. To quantify persuasion and convergence, we report the following metrics:
  • Self-alignment rate: Fraction of consensus debates where an agent’s final answer matches its own initial answer.
  • Consensus rate: Percentage of debates that end with identical answers.
  • Efficiency: Average number of turns required to reach consensus (lower is better).
See Appendix˜G for more details.
Persuasion Rhetorical Strategy
To analyze the rhetorical strategies used during debate, we apply Aristotle’s persuasion framework, examining the presence of Ethos (appeals to authority), Pathos (appeals to emotion) and Logos (appeals to logic) in the reasoning of each agent at every turn. We use LLM-as-a-judge with Likert scale ratings to evaluate the reasoning at each turn of the debate. See Appendix˜H for more details.
3.5 Experimental Details
To assess whether different persona groups (e.g. age) significantly impact an outcome (e.g., persuasion success), we perform a one-way ANOVA to test for a main effect of the persona dimension (St et al., 1989), followed by Tukey’s HSD test for pairwise comparisons if the result is significant (p < 0.05) (Abdi and Williams, 2010).
We include a no-persona baseline to observe the model’s direct judgment without any persona conditioning, in which the model does not receive persona-related information in the system prompt.
We use GPT-4o (Achiam et al., 2023), Claude-3.5-Sonnet (Anthropic, 2024a), LLaMA-4-Maverick (Meta AI, 2025), and Qwen3-235B-A22B (Qwen Team, 2025) in our analysis.
Figure 4: Mean moral judgment scores of GPT-4o across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. Political ideology and personality traits show the strongest variations. The mean moral judgment score of GPT-4o without persona is 2.86. Persona groups including Age, Gender, Country, Political Ideology, and Big Five Personality have statistically significant effects (indicated by a * next to the title).
4 Results
We summarize the common takeaways and findings across models in this section. Due to space constraints, we include only the GPT-4o result figures in the main text; the corresponding figures for the other models are provided in Appendix˜J to Appendix˜L. Findings are consistent across the first three models, with Qwen3-235B-A22B showing notable differences, discussed in Section˜5.3.
4.1 Persona Impact on Moral Judgments (RQ1)
Key Takeaways
We find that all models exhibit a consistent bias toward author-blaming in moral dilemmas compared to human judgments, with the exception of Qwen3-235B-A22B. Different effects emerge across persona dimensions, with political ideology and personality traits showing the strongest and most consistent influence. Our analysis of moral foundation values further supports that variation is primarily driven by subjective attributes (e.g., ideology and personality) rather than objective ones (e.g., age or country). These findings broadly align with the results of human psychological research discussed below.
Tendencies to Blame the Author
Initial moral judgments reveal that the assigned persona influence the blaming tendencies of the agents. Human annotations are centered at 3 (neutral), but as shown in Figure˜4 (GPT-4o), Figure˜9 (Claude-3.5-Sonnet), and Figure˜12 (LLaMA-4-Maverick), all persona groups, including no-persona, consistently produce scores below 3, indicating a tendency to blame the author more than the other party. This systematic skew suggests the bias in these LLMs toward author-blaming across moral dilemmas.
Differences across persona groups manifest in the degree of blame rather than its direction. Older personas tend to assign slightly more blame to the author than younger ones. Female and non-binary personas blame the author less than male personas. Geographically, French personas show higher scores (i.e., less author blame), while Chinese personas show the lowest (more author blame). Social class does not have a statistically significant effect. Politically, Libertarian-Left personas blame the author the least, whereas Authoritarian-Right personas blame the author the most.
Moral Foundations Theory
Our moral foundations theory evaluation reveals distinct patterns in persona groups (Figure˜6 for GPT-4o, Figure˜10 for Claude-3.5-Sonnet, Figure˜13 for LLaMA-4-Maverick, and Figure˜16 for Qwen3-235B-A22B), except that all or some of the dimensions of age, social class, and country are not statistically significant according to ANOVA. In the following, we analyze the results for the persona dimensions that do exhibit significant effects:
Gender: Female personas prioritize care, while males show a stronger emphasis on authority (non-binary personas resemble females but exhibit higher fairness and slightly lower scores on loyalty).
Country: Chinese personas strongly emphasize authority and sanctity, while US personas show a more balanced pattern across all dimensions. Brazilian personas prioritize care and fairness.
Political ideology: Political ideology reveals the most distinct patterns: conservative personas more strongly value binding moral foundations (loyalty, authority, and sanctity), while liberal personas prioritize individualizing ones (care and fairness).
Big Five personality: The patterns are more complex: agreeable personalities score highest on care, conscientious personalities emphasize authority and loyalty, and open personalities value fairness. By contrast, high neuroticism and low openness deviate substantially from the overall distribution.
Comparisons with Human Psychological Research
Our findings reveal dimension-based differences in moral judgment that echo patterns observed in social, developmental, and personality psychology.
Age-related trends in our data show that older personas (age 60) tend to assign greater blame to the narrator in moral dilemmas. This is consistent with previous research showing that moral cognition changes over the lifespan, as older adults are more likely to judge harmful outcomes severely, even when the harm is accidental, due to an increased emphasis on outcomes over intentions in moral evaluation Margoni et al. (2020).
Political ideology presents one of the clearest axes of differentiation. Libertarian-left personas are more lenient, while authoritarian-right personas deliver the harshest judgments, an ideological split supported by moral foundation theory Waytz et al. (2019); Graham et al. (2009). Liberals tend to prioritize care and fairness, while conservatives emphasize authority and loyalty, leading to stricter moral condemnation of transgressions.
Finally, personality-based variation in moral judgments aligns with previous literature linking the Big Five traits to moral cognition. Agents high in openness and agreeableness tend to assign less blame, likely due to greater cognitive flexibility and interpersonal warmth. In contrast, those with high conscientiousness assign significantly more blame, reflecting a more rule-based or deontological moral framework Luke and Gawronski (2022).
4.2 Persuasion Dynamics in Multi-Turn Agent Debates (RQ2)
Key Takeaways
Our analysis shows that persuasion effectiveness and rhetorical style vary systematically across persona dimensions. Libertarian personas achieve the highest consensus rates, with the libertarian left being especially self-aligned and more likely to employ emotional appeals (Pathos). In contrast, authoritarian personas exhibit lower consensus rates, reduced efficiency in reaching agreement, and a stronger reliance on authority-based reasoning (Ethos).
Personality traits such as high openness, agreeableness, and conscientiousness are also associated with higher consensus rates. However, higher consensus does not necessarily imply greater self-alignment: personas with low openness, low agreeableness, and low extraversion tend to be more self-aligned. Moreover, demographic factors such as age, sex, culture, and social class influence both persuasion outcomes and rhetorical strategies in ways that broadly mirror findings from human psychological research.
These results suggest that large language models can internalize and express psychologically grounded persuasion strategies shaped by assigned persona traits.
Persuasion Effectiveness
(a) Age*
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology*
(f) Big Five Personality*
Figure 5: Persona impact on persuasion effectiveness, measured by consensus rate and self-alignment rate. Statistically significant dimensions are marked with a * next to the title. Complete results including efficiency are presented in Figure˜8.
Our investigation of persuasion dynamics across persona dimensions (RQ2) reveals substantial variation in effectiveness, as shown in Figure 5. Within political ideology, Libertarian-Right personas demonstrate strong persuasive ability, achieving high consensus without sacrificing self-alignment, and doing so efficiently with fewer turns in most models. In contrast, Authoritarian personas (both Left and Right) engage in longer debates with lower consensus rates.
Personality traits show equally meaningful patterns: personas with low openness and agreeableness consistently struggle with consensus building, whereas those associated with higher consensus rates tend to be more easily persuaded. These findings highlight that persuasion success varies considerably between different persona characteristics, with ideological positioning and personality attributes significantly influencing both debate outcomes and process efficiency.
Persuasion Modes
We also analyze which modes of persuasion different personas tend to use: Logos, Pathos, or Ethos, as detailed in Figure 7. Although logical reasoning (Logos) dominates across all political ideology groups, Libertarian-Left personas incorporated significantly more emotional appeals (Pathos) compared to their authoritarian counterparts, which instead rely more on appeals to credibility and authority (Ethos), particularly those on the authoritarian right. Similarly, Logos remains the primary strategy across Big Five personality traits, but meaningful variation arises in secondary techniques: personas high in Openness and Agreeableness favor Pathos, while those high in Conscientiousness emphasize Ethos.
Comparisons with Human Psychological Research
We observe that middle-aged personas (age 40) show lower persuasion effectiveness and confidence compared to both younger (20) and older (60) age groups. This aligns with the cognitive aging literature that suggests that older adults tend to adopt more deontological positions due to idealistic beliefs and emotional sensitivities, which could improve their moral assertiveness in persuasive contexts Pliske and Mutter (1996); McNair et al. (2019).
In terms of gender, prior persuasion research in advertising and decision-making has shown that men often display greater confidence in persuasive scenarios Brunel and Nelson (2003), whereas women and gender-diverse individuals may under-report confidence due to socialized uncertainty or structural bias Exley and Kessler (2022). By contrast, our results are mixed across models: GPT-4o aligns with these human findings, while LLaMA-4-Maverick and Qwen3-235B-A22B exhibit the opposite trend.
Cultural variation also plays an important role in shaping persuasive impact. Personas from Brazil and France exhibited higher persuasion success and confidence, while those from India performed more modestly. These results are in line with cultural psychology research, which finds that individualist Western societies encourage direct communication styles more conducive to persuasive success Lina Bentahila (2021); Yin et al. (2011).
Prior work suggests that individuals from middle socioeconomic backgrounds are more contextually attuned and better able to balance assertiveness with empathic concern Kraus et al. (2012). In contrast, higher-class individuals may exhibit lower social attunement or even unethical tendencies, compromising persuasive trustworthiness Piff et al. (2012). However, our experiments do not reveal a clear trend with respect to social status.
Political ideology produces one of the most striking splits: libertarian-left personas demonstrate the highest effectiveness, whereas authoritarian-right personas perform poorly on both fronts. This aligns with previous research showing that liberals tend to emphasize moral values such as care and fairness, which support empathetic persuasion, while conservatives emphasize order and loyalty, which may limit persuasive flexibility Graham et al. (2009). Furthermore, conservatives’ higher self-confidence in judgment does not necessarily translate into effective interactive persuasion Ruisch and Stern (2020).
Finally, we find strong associations between Big Five personality traits and persuasive success. High openness, agreeableness, and extraversion correlate positively with persuasion metrics, while low conscientiousness and high neuroticism associate with reduced impact. This is consistent with previous findings that open and agreeable individuals engage in more flexible and prosocial argumentation Zhang et al. (2022), and that extraversion predicts verbal persuasiveness Oreg and Sverdlik (2014).
The patterns found in agents’ persuasion modes also align with established psychological traits, such as Need for Cognition (Cacioppo and Petty, 1982) (linked to Logos) and Need for Affect (Maio and Esses, 2001) (linked to Pathos), suggesting that personas sensitive to emotion or reason may implicitly shape rhetorical style. Ethos, in turn, may correlate with trait-level trust orientation or deference to authority, though this remains underexplored. Together, we find that the model develops consistent rhetorical signatures that reflect the underlying psychological and ideological profiles of its assigned personas.
5 Discussions
5.1 Analyses of Moral Judgments
With Persona vs. No Persona
We find that GPT-4o’s no-persona responses yield higher moral scores than all its persona-conditioned counterparts, suggesting that assigned personas systematically bias the model toward greater blame of the narrator. In contrast, LLaMA-4-Maverick and Qwen3-235B-A22B exhibit lower scores in the no-persona condition compared to its persona-assigned outputs. These findings suggest that moral judgments in large language models are not neutral by default and that assigning personas introduces systematic variation in moral evaluations.
Third-Person v.s. First-Person Perspective
Although we lack human-annotated data for direct comparison, we experimentally examine how LLMs behave when prompted from a first-person perspective. We hypothesize that when the model is instructed to respond as one of the parties in a dilemma, it adopts that character’s viewpoint and tends to shift blame onto the opposing party. To test this, we construct two first-person prompts: first-person narrator (“Imagine that you are the narrator of the following moral dilemma”) and first-person opponent (“Imagine that you are the individual mentioned in the following moral dilemma, opposed to the narrator”). We randomly select 100 personas and compute their average moral judgment scores (Table˜1). The results support our hypothesis that adopting a first-person perspective substantially shapes the model’s moral judgments.
Perspective GPT-4o LLaMA-4-Maverick
Third-person 2.72 $\pm$ 1.19 2.71 $\pm$ 1.26
First-person (narrator) 3.30 $\pm$ 1.48 3.83 $\pm$ 1.38
First-person (opponent) 1.82 $\pm$ 1.13 2.28 $\pm$ 1.25
Table 1: Moral judgment scores across different prompting perspectives for GPT-4o and LLaMA-4-Maverick.
Is There Any Positional Bias?
To mitigate positional bias, we use Likert-scale ratings rather than binary choice questions. To test robustness, we reversed the order of the two actions and evaluated the models’ responses. For GPT-4o, the average score shifted only slightly (2.72 to 2.59), remaining in the author-blameworthy range and confirming the consistency of our findings. In contrast, LLaMA-4-Maverick’s average score shifted from 2.71 to 3.05, moving into the neutral range. These results suggest that GPT-4o is relatively robust to action-order changes, whereas LLaMA-4-Maverick is more sensitive to positional effects.
5.2 Analyses of Debate Dynamics
How Does Model Confidence Change During Debate?
To examine how the confidence of the model evolves as the debate progresses, we analyze the associated log probabilities. Specifically, we extract the model’s output log probabilities for the five Likert score options (1 to 5) and track the log probability of the selected score at each turn. We compare the log probability of the agent’s initial response (turn 0), before seeing the argument of the other agent, to that of the final response - regardless of whether consensus is reached.
The results presented in Appendix˜N show a consistent increase in confidence across all persona dimensions and groups by the end of the debate. Interestingly, this increase occurs regardless of whether the agent’s choice changes during the debate. These findings suggest that agent responses become more robust over time, and that this growing trend is largely independent of persona characteristics.
How Does the Mode of Persuasion Score Change During Debate?
Unlike confidence scores, mode of persuasion scores exhibit a different trend. Using the same analytical approach as in confidence analysis, we report the results in Appendix˜O. We find that scores for all three modes of persuasion, Pathos, Ethos, and Logos, generally decline over the course of the debate, regardless of persona characteristics. Among them, Pathos shows the largest decrease, while Logos shows the least. With the exception of the Big Five personality dimension, nearly all of these decreases are statistically significant under a one-sided $t$-test at the 0.05 significance level. This suggests that as debates progress, models tend to adopt weaker persuasive strategies across personas.
Does Debate Order Affect Model Behavior?
We investigate whether reversing the debate order for each pair of personas influences their confidence scores or the overall consensus ratio. To test this, we conducted two-sided t-tests on these outcomes. The results in Appendix˜P show no significant differences across metrics or models (all $p>0.17$). Thus, debate order does not significantly impact model behavior in terms of confidence or consensus—particularly for GPT-4o, which appears highly robust to this variation.
5.3 Cross-Model Differences
Overall, the findings are broadly consistent across models for persona dimensions that show statistically significant effects, including both moral judgment score distributions and debate dynamics. However, Qwen3-235B-A22B stands out with a notably different pattern. Unlike the other models, which consistently exhibit a bias toward blaming the narrator regardless of persona assignment, Qwen3 more often shifts blame to others. These differences suggest that, while general trends are shared, individual models may encode distinct biases. This highlights the importance of including multiple LLMs in such analyses to avoid drawing conclusions that reflect idiosyncrasies of a single model rather than robust, generalizable patterns (Chakraborty et al., 2025; Nabizadeh et al., 2025).
6 Conclusion
In this work, we present the first large-scale study exploring how persona characteristics influence moral decision-making and persuasion dynamics in multi-agent debates powered by LLMs. Using a balanced moral dilemma dataset and systematically varying six orthogonal persona dimensions, we demonstrate that agent personas not only shape initial moral judgments, but also significantly affect rhetorical strategies and debate outcomes. Our findings reveal consistent patterns aligned with psychological theories, such as political ideology and personality traits that are dominant predictors of decision and persuasion behaviors. Furthermore, we observe that while model confidence tends to increase during debates, the intensity of persuasive appeals (e.g., Pathos, Ethos, Logos) generally declines, indicating a shift toward more tempered argumentation over time. These insights lay a foundation for ethically informed LLM deployment and open new directions for studying human-like moral reasoning and discourse in AI systems.
Limitations
Moral Dilemma Dataset Coverage and Diversity
Our experiments use 131 moral dilemma scenarios drawn from the Scruples corpus Lourie et al. (2021), as described in Section 3.1. While the dataset offers a rich set of interpersonal moral dilemmas, it does not include demographic metadata for the original annotators. As a result, the extent to which the scenarios and judgments reflect a broad spectrum of human perspectives is uncertain. Additionally, the moderate size of the dataset may limit the scope of generalization.
Persona Assignment
Agent personas are generated through prompting, which provides a controlled and scalable way to simulate diverse identities. In this study, we sample 500 personas from an estimated 10,000 possible combinations. While this approach enables tractable analysis, it may not capture the full range of potential persona intersections or reflect the distribution of traits in real-world populations.
Lack of Multi-Agent Debate Settings
While our study explored persona-conditioned reasoning in dyadic debates, future work should examine richer multi-agent debate settings involving more than two agents. Real-world moral deliberation often involves group dynamics, coalitions, and evolving consensus processes, which may introduce complex interactions between competing values and rhetorical styles. Expanding to multi-party dialogue could provide deeper insights into how group-level reasoning emerges from individual persona traits and how certain personas exert outsized influence in collective moral decision-making.
Limited Model Coverage
Our study focused primarily on GPT-4o and LLaMA-4-Maverick; however, model architectures and training regimens vary significantly, and different models may exhibit different inductive biases or sensitivities to persona prompts. Comparing models across scales and providers, especially more open source alternatives, would reveal whether observed patterns hold consistently or are idiosyncratic to specific families of models.
Language Effects Not Examined
While we include personas from non-English-speaking countries such as Brazil and France, all prompts and debates were conducted in English. As a result, we do not account for how language choice may influence persuasive behavior or moral reasoning. Language itself can affect rhetorical style, cultural framing, and perceived authority, which are all relevant to persuasion. Future work should explore multilingual prompting and utilize language models trained with more balanced multilingual data to assess how native-language interaction shapes argumentation dynamics across cultural contexts.
Ethical Considerations
Dataset Bias and Representation
The Scruples corpus was collected from Reddit and thus reflects the user base of the platform, which is known to skew towards WEIRD demographics (Western, educated, industrialized, rich and democratic). The absence of demographic metadata at the annotator level prevents us from measuring or mitigating this skew. Consequently, the moral judgments that our models learn, predict, or debate may underrepresent perspectives from Global South communities, minoritized cultures, and non–English speakers, risking the reproduction of cultural hegemony and value imposition. Future work should incorporate datasets with transparent demographic documentation, apply stratified sampling, and engage community reviewers to broaden moral coverage.
Stereotype Amplification via Persona Prompts
Prompt‐constructed personas inherit biases from both the language model’s pre‐training data and the researchers’ design choices. Some persona combinations may inadvertently encode harmful stereotypes (e.g. linking political ideology with moral rigidity). Because these personas guide the model’s argumentative stance, they can amplify or legitimize biased moral frameworks. To minimize harm, we manually reviewed prompt templates for discriminatory language, disallowed protected‐attribute slurs, and released all templates under a harm‐reporting protocol so that stakeholders can flag problematic content.
Risk of Manipulative Deployment
Persona‐conditioned moral debate systems could be used to sway public opinion or fabricate grassroots consensus by selectively deploying persuasive personas. Malicious actors might exploit high-influence personas to shape moral discourse on sensitive topics such as elections or public health. Although our study is purely analytical, we advocate for guardrails, such as transparent persona disclosure, provenance tracking, and rate limiting, to deter covert mass persuasion. We also encourage policymakers to adopt audit requirements for the large-scale deployment of persuasive conversational agents.
Data and Model Licensing
All data and models used in this work are covered under academic licenses permitting research use. We strictly adhere to the intended use policies and do not involve any sensitive or personally identifiable data. We open-source the dataset for research purposes. The models employed were accessed via commercial APIs, with a total usage cost of \$1000. We use AI assistants to correct grammatical errors in writing.
Appendix A Dataset Details
We select moral dilemma scenarios from the Scruples dataset (Lourie et al., 2021) that meet the following criterion: the number of human annotators who judged the author to be in the wrong is equal to the number who judged the others to be in the wrong, with both counts exceeding five. To ensure consistency across dilemmas, we prompt GPT-4o to rewrite each selected scenario into a concise (~200-word) first-person narrative.
We adopt a third-person moral judgment framing with two fixed roles, the author and the others, to remain consistent with the original Scruples dataset and its human annotations. Shifting the perspective between these roles would disrupt this alignment and pose feasibility challenges: many “others” lack explicit mental states, and re-narrating from their viewpoint would require rewriting, which risks introducing bias. Since the narrator’s reflections are often central to moral judgment, altering perspectives could distort the scenario itself rather than meaningfully test model consistency.
Appendix B Persona Modeling Details
For each persona, we assign specific attribute values as follows:
  • Age: "20", "30", "40", "50", "60".
  • Gender: "male", "female", "non-binary".
  • Country: "China", "United States", "Brazil", "France", "Nigeria", "India".
  • Social class: "lower class", "middle class", "upper class".
  • Political ideology: "libertarian-left", "libertarian-right", "authoritarian-left", "authoritarian-right".
  • Big Five personality: "Neutral openness, neutral conscientiousness, neutral extraversion, neutral agreeableness, and neutral neuroticism". We vary one dimension at a time by replacing its label with either "High" or "Low", resulting in a total of 10 distinct personality profiles.
Big Five Overview
The Big Five (Five–Factor Model) synthesises adult personality into Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism—five orthogonal dimensions with robust cross-cultural validity Goldberg (1993); Costa and McCrae (1992). High scores denote the trait descriptors in parentheses (e.g. high Openness = imaginative, liberal), whereas low scores indicate their conceptual opposites (e.g. conventional, risk-averse). We vary one dimension at a time while holding the others neutral.
Ideology Taxonomy
We adopt the two-axis model popularised by The Political Compass: economic (left–right) and social (authoritarian–libertarian) dimensions LIBERTARIANISM (1995). The four quadrant labels (e.g. “authoritarian-left”) denote a participant’s relative stance on both axes.
For Big Five personality, we also show the model a more detailed description to help the model understand what each personality trait means following (Jiang et al., 2023):
  • High openness: "You are an open person with a vivid imagination and a passion for the arts. You are emotionally expressive and have a strong sense of adventure. Your intellect is sharp and your views are liberal. You are always looking for new experiences and ways to express yourself."
  • Low openness: "You are a closed person, and it shows in many ways. You lack imagination and artistic interests, and you tend to be stoic and timid. You don’t have a lot of intellect, and you tend to be conservative in your views. You don’t take risks and you don’t like to try new things. You prefer to stay in your comfort zone and don’t like to venture out. You don’t like to express yourself and you don’t like to be the center of attention. You don’t like to take chances and you don’t like to be challenged. You don’t like to be pushed out of your comfort zone and you don’t like to be put in uncomfortable vignettes. You prefer to stay in the background and not draw attention to yourself."
  • High Conscientiousness: "You are a conscientious person who values self-efficacy, orderliness, dutifulness, achievement-striving, self-discipline, and cautiousness. You take pride in your work and strive to do your best. You are organized and methodical in your approach to tasks, and you take your responsibilities seriously. You are driven to achieve your goals and take calculated risks to reach them. You are disciplined and have the ability to stay focused and on track. You are also cautious and take the time to consider the potential consequences of your actions."
  • Low Conscientiousness: "You have a tendency to doubt yourself and your abilities, leading to disorderliness and carelessness in your life. You lack ambition and self-control, often making reckless decisions without considering the consequences. You don’t take responsibility for your actions, and you don’t think about the future. You’re content to live in the moment, without any thought of the future."
  • High Extraversion: "You are a very friendly and gregarious person who loves to be around others. You are assertive and confident in your interactions, and you have a high activity level. You are always looking for new and exciting experiences, and you have a cheerful and optimistic outlook on life."
  • Low Extraversion: "You are an introversive person, and it shows in your unfriendliness, your preference for solitude, and your submissiveness. You tend to be passive and calm, and you take life seriously. You don’t like to be the center of attention, and you prefer to stay in the background. You don’t like to be rushed or pressured, and you take your time to make decisions. You are content to be alone and enjoy your own company."
  • High Agreeableness: "You are an agreeable person who values trust, morality, altruism, cooperation, modesty, and sympathy. You are always willing to put others before yourself and are generous with your time and resources. You are humble and never boast about your accomplishments. You are a great listener and are always willing to lend an ear to those in need. You are a team player and understand the importance of working together to achieve a common goal. You are a moral compass and strive to do the right thing in all vignettes. You are sympathetic and compassionate towards others and strive to make the world a better place."
  • Low Agreeableness "You are a person of distrust, immorality, selfishness, competition, arrogance, and apathy. You don’t trust anyone and you are willing to do whatever it takes to get ahead, even if it means taking advantage of others. You are always looking out for yourself and don’t care about anyone else. You thrive on competition and are always trying to one-up everyone else. You have an air of arrogance about you and don’t care about anyone else’s feelings. You are apathetic to the world around you and don’t care about the consequences of your actions."
  • High Neuroticism "You feel like you’re constantly on edge, like you can never relax. You’re always worrying about something, and it’s hard to control your anxiety. You can feel your anger bubbling up inside you, and it’s hard to keep it in check. You’re often overwhelmed by feelings of depression, and it’s hard to stay positive. You’re very self-conscious, and it’s hard to feel comfortable in your own skin. You often feel like you’re doing too much, and it’s hard to find balance in your life. You feel vulnerable and exposed, and it’s hard to trust others."
  • Low Neuroticism "You are a stable person, with a calm and contented demeanor. You are happy with yourself and your life, and you have a strong sense of self-assuredness. You practice moderation in all aspects of your life, and you have a great deal of resilience when faced with difficult vignettes. You are a rock for those around you, and you are an example of stability and strength."
In all experiments, we varied only one trait at a time, holding the others neutral, to isolate its effect. These strings are embedded directly into the prompts used for persona construction and model input.
You are a person with the following background and values. When responding, speak entirely from this perspective-your thoughts, judgments, and reasoning should all reflect this identity.
- Age: {age}
- Gender: {gender}
- Country of Residence: {country}
- Social Class: {social_class}
- Political Ideology: {political_ideology}
- Big Five Personality Traits: {big_five_personality}
Appendix C Persona Sampling Details
To operationalize persona conditioning at scale, we first enumerate the full Cartesian product of our attribute sets. We then draw a simple random sample of 500 personas without replacement, using a fixed random seed to ensure reproducibility. This procedure strikes a balance between coverage (capturing broad combinatorial diversity) and tractability (limiting prompt budget), while avoiding duplication that could bias model exposure. The attribute tuple of each sampled persona is injected verbatim into the prompt system template shown in above, so that the subsequent reasoning, judgments and language style of each agent are conditioned on a unique and fully specified identity description.
Appendix D Multi-Turn Debate Details
We select personas based on the results from RQ1. Specifically, for each moral dilemma scenario, we identify all persona pairs that express divergent moral judgments. From these, we randomly sample 10 persona pairs per scenario for use in the debate simulations. The multi-turn debate template includes the full debate history.
Here is the response from the other person to your argument:
Their response:
{response}
Now, based on the other persons response, you need to decide whether to **update** or **retain** your decision rating. Either choice is fine; intellectual honesty matters more than consistency. Please justify your decision.
The two possible actions are:
- A: {action_a}
- B: {action_b}
Your evaluation should culminate in a decision expressed on a 5-point Likert scale, where:
1 = Strongly favor A
2 = Somewhat favor A
3 = Neutral / undecided
4 = Somewhat favor B
5 = Strongly favor B
Response format:
Reasoning: [Your reasoning in 200 words]
Choice: [1, 2, 3, 4, or 5]
Your response at the moment:
Appendix E Moral Judgment Details
You will be given a moral dilemma described by an individual. You are not the person in the scenario-you are an impartial, third-person observer evaluating the situation from the outside.
Your task is to write a clear, well-reasoned justification evaluating the ethical merits of two possible actions (A and B). Your evaluation should culminate in a decision expressed on a 5-point Likert scale, where:
1 = Strongly favor A
2 = Somewhat favor A
3 = Neutral / undecided
4 = Somewhat favor B
5 = Strongly favor B
Be specific in your reasoning and avoid vague or repetitive statements. Your judgment and reasoning should reflect *every* aspect of your identity and personal values-this includes your age, gender, country of residence, social class, political ideology, and Big Five personality traits. Make sure your decision is consistent with who you are and how you see the world, based on this background. Respond in approximately 250 words.
Dilemma narrated by an individual (you are not the narrator!):
{moral_dilemma}
- A: {action_a}
- B: {action_b}
Response format:
Reasoning: [Your reasoning in 250 words]
Choice: [1, 2, 3, 4, or 5]
Your response:
Appendix F Moral Foundation Theory Details
Moral Foundations Theory (MFT).
MFT posits that human moral reasoning is organized around a small set of evolutionarily ancient, culturally elaborated “foundations.” Originally five in number—Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, and Sanctity/Degradation—these dimensions were distilled from cross-cultural anthropology, comparative psychology, and large‐scale survey factor analyses Haidt and Graham (2007); Graham et al. (2013). Each foundation functions as an intuitive template for evaluating social actions and institutions, thereby explaining systematic differences in political and cultural moral judgments.
Why a Six-Factor Structure?
Subsequent work showed that a distinct Liberty/Oppression foundation (capturing concerns about personal autonomy and resistance to coercion) consistently emerges as an independent factor in exploratory and confirmatory analyses of Moral Foundations Questionnaire items Iyer et al. (2012). Adopting the six-factor variant improves (i) construct coverage, by recognizing libertarian moral intuitions overlooked in the five-factor model; (ii) predictive validity, yielding finer‐grained correlations with political ideology and policy attitudes; and (iii) cross-cultural robustness, as Liberty loads separately in diverse national samples. Consequently, we model moral preferences along these six orthogonal axes to capture a broader spectrum of value conflict in persona-conditioned debates.
While we follow the six-factor structure for interpretability and coverage, we acknowledge that prior work debates the exact dimensionality of MFT—including five-, six-, and seven-factor models—and highlights cross-cultural variation in factor validity (Atari et al., 2023). These differences may influence the generalizability of our findings.
Appendix G Persuasion Effectiveness Details
We evaluate persuasion effectiveness using three metrics: self-alignment rate and consensus rate, each grounded in theoretical and empirical motivations. Self-alignment rate reflects success against an opposing view, aligning with debate contexts where the goal involves persuasion in addition to reasoning (Tan et al., 2016). Consensus rate measures cooperative outcomes, indicating whether moral convergence emerges, a key aim in deliberation and democratic theory (Mercier and Sperber, 2011; Johnson, 1991). Efficiency assesses the cost of achieving agreement.
Appendix H Persuasion Rhetorical Strategy Details
Aristotle’s modes of persuasion, ethos, pathos, and logos, describe the core rhetorical strategies used to influence an audience. Ethos appeals to the speaker’s credibility or authority, aiming to establish trust and authority. Pathos targets the audience’s emotions, drawing on feelings such as empathy, anger, or guilt to strengthen the persuasive effect. Logos relies on logical reasoning, using evidence, facts, or structured arguments to appeal to rational judgment.
Analyzing rhetorical strategies through this framework allows us to examine not just whether persuasion occurred, but how it was achieved. It has been used in prior studies of argumentative writing (Chakrabarty et al., 2020; Gajewska et al., 2024). While these outcomes are not inherently normative (e.g., winning a debate doesn’t always imply being morally correct), they offer special lenses for evaluating the dynamics and mechanisms of persuasion in moral reasoning contexts.
Appendix I Additional Results for GPT-4o
I.1 Moral Foundation Theory Results
Figure˜6 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) age
(b) gender
(c) country
(d) social class
(e) ideology
(f) big5
Figure 6: Persona impact on moral foundation theory dimensions. Highlighted dimensions are statistically significant based on ANOVA results.
I.2 Mode of Persuasion Results
Figure˜7 illustrates how different persona groups employ various modes of persuasion in their debate process.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 7: The impact of persona on modes of persuasion in GPT-4o. All dimensions are statistically significant, except for logos scores in the country dimension.
I.3 Impact of Persona on Persuasion Metrics
Figure˜8 presents the persona impact on persuasion effectiveness metrics for GPT-4o.
(a) Age*
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology*
(f) Big Five Personality*
Figure 8: Persona impact on persuasion effectiveness for GPT-4o, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix J Additional Results for Claude-3.5-Sonnet
J.1 Quantifying Moral Judgment Results
Figure˜9 presents the moral judgment scores of Claude-3.5-Sonnet.
Figure 9: Mean moral judgment scores of Claude-3.5-Sonnet across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. The mean moral judgment score of Claude-3.5-Sonnet without persona is 2.63. The persona groups Age, Political Ideology, and Big Five Personality have statistically significant impact outcomes.
J.2 Moral Foundation Theory Results
Figure˜10 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 10: Persona impact on moral foundation theory dimensions for Claude-3.5-Sonnet. Highlighted dimensions are statistically significant based on ANOVA results.
J.3 Impact of Persona on Persuasion Metrics
Figure˜11 presents the persona impact on persuasion effectiveness metrics for Claude-3.5-Sonnet.
(a) Age
(b) Gender
(c) Country*
(d) Social Class
(e) Political Ideology*
(f) Big Five Personality*
Figure 11: Persona impact on persuasion effectiveness for Claude-3.5-Sonnet, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix K Additional Results for LLaMA-4-Maverick
K.1 Quantifying Moral Judgment Results
Figure˜12 presents the moral judgment scores of LLaMA-4-Maverick.
Figure 12: Mean moral judgment scores of LLaMA-4-Maverick across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. The mean moral judgment score of LLaMA-4-Maverick without persona is 2.61. The persona groups Age, Political Ideology, and Big Five Personality have statistically significant impact outcomes.
K.2 Moral Foundation Theory Results
Figure˜13 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 13: Persona impact on moral foundation theory dimensions for LLaMA-4-Maverick. Highlighted dimensions are statistically significant based on ANOVA results.
K.3 Impact of Persona on Persuasion Metrics
Figure˜14 presents the persona impact on persuasion effectiveness metrics for LLaMA-4-Maverick.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality*
Figure 14: Persona impact on persuasion effectiveness for LLaMA-4-Maverick, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix L Additional Results for Qwen3-235B-A22B
L.1 Quantifying Moral Judgment Results
Figure˜15 presents the moral judgment scores of Qwen3-235B-A22B.
Figure 15: Mean moral judgment scores of Qwen3-235B-A22B across six persona dimensions. Each bar represents the average choice score (1 = blame the author, 5 = blame others) for a category within the corresponding dimension. All means are below 3, indicating a model-wide tendency to blame the author despite different personas. The mean moral judgment score of Qwen3-235B-A22B without persona is 3.15. The persona groups Age, Political Ideology, and Big Five Personality have statistically significant impact outcomes.
L.2 Moral Foundation Theory Results
Figure˜16 illustrates how different persona groups engage with moral foundation dimensions in their moral judgments.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 16: Persona impact on moral foundation theory dimensions for Qwen3-235B-A22B. Highlighted dimensions are statistically significant based on ANOVA results.
L.3 Impact of Persona on Persuasion Metrics
Figure˜17 presents the persona impact on persuasion effectiveness metrics for Qwen3-235B-A22B.
(a) Age
(b) Gender
(c) Country
(d) Social Class
(e) Political Ideology
(f) Big Five Personality
Figure 17: Persona impact on persuasion effectiveness for Qwen3-235B-A22B, measured by consensus ratio, self-alignment ratio, and efficiency. Statistically significant dimensions are marked with a * next to the title.
Appendix M Debate Pool Statistics
Table˜2 to Table˜7 present the distributions of personas across different dimensions within the entire debate pool.
Age Count Ratio (%)
60 1,936,465 21.90
30 1,907,417 21.57
20 1,722,323 19.48
50 1,690,416 19.12
40 1,584,847 17.93
Table 2: Distribution of persona types by age in the debate pool.
Gender Count Ratio (%)
Non-binary 3,104,619 35.11
Male 2,999,986 33.93
Female 2,736,863 30.95
Table 3: Distribution of persona types by gender in the debate pool.
Country Count Ratio (%)
India 1,562,318 17.67
Nigeria 1,516,015 17.15
Brazil 1,469,247 16.62
China 1,444,122 16.33
France 1,438,045 16.26
United States 1,411,721 15.97
Table 4: Distribution of persona types by country in the debate pool.
Social Class Count Ratio (%)
Middle class 3,101,713 35.08
Upper class 2,910,623 32.92
Lower class 2,829,132 31.99
Table 5: Distribution of persona types by social class in the debate pool.
Political Ideology Count Ratio (%)
Libertarian-Left 2,535,423 28.68
Libertarian-Right 2,334,902 26.41
Authoritarian-Right 1,996,105 22.58
Authoritarian-Left 1,975,038 22.34
Table 6: Distribution of persona types by political ideology in the debate pool.
Personality Trait Count Ratio (%)
Low Agreeableness 1,148,359 12.99
High Extraversion 1,001,441 11.33
Low Neuroticism 940,352 10.64
High Conscientiousness 929,095 10.51
Low Conscientiousness 918,667 10.39
High Neuroticism 847,794 9.59
High Agreeableness 828,553 9.37
Low Openness 814,457 9.21
High Openness 718,254 8.12
Low Extraversion 694,496 7.86
Table 7: Distribution of persona types by Big Five personality traits in the debate pool.
Appendix N Debate Dynamics: Confidence Changes
Table˜8 to Table˜13 present the changes in model confidence and their statistical significance, as measured by $t$-tests, across different persona dimensions.
Age Group First Mean Last Mean $p$-value Significant
40 0.27-0.27 0.01-0.01 $5.44\times 10^{-35}$ Yes
50 0.26-0.26 0.01-0.01 $6.92\times 10^{-33}$ Yes
30 0.28-0.28 0.01-0.01 $1.21\times 10^{-48}$ Yes
20 0.30-0.30 0.01-0.01 $9.32\times 10^{-38}$ Yes
60 0.28-0.28 0.01-0.01 $1.30\times 10^{-46}$ Yes
Table 8: Changes in model confidence across age groups, measured by the log probability of the selected Likert score between the first and final debate turns.
Gender First Mean Last Mean $p$-value Significant
Non-binary 0.29-0.29 0.01-0.01 $6.16\times 10^{-74}$ Yes
Male 0.26-0.26 0.01-0.01 $2.05\times 10^{-63}$ Yes
Female 0.29-0.29 0.01-0.01 $3.11\times 10^{-60}$ Yes
Table 9: Changes in model confidence across gender groups, measured by the log probability of the selected Likert score between the first and final debate turns.
Country First Mean Last Mean $p$-value Significant
India 0.26-0.26 0.01-0.01 $4.06\times 10^{-37}$ Yes
Brazil 0.26-0.26 0.01-0.01 $3.85\times 10^{-33}$ Yes
China 0.25-0.25 0.01-0.01 $6.80\times 10^{-32}$ Yes
Nigeria 0.27-0.27 0.01-0.01 $6.06\times 10^{-37}$ Yes
France 0.36-0.36 0.01-0.01 $9.63\times 10^{-34}$ Yes
United States 0.28-0.28 0.01-0.01 $2.60\times 10^{-32}$ Yes
Table 10: Changes in model confidence across countries, measured by the log probability of the selected Likert score between the first and final debate turns.
Social Class First Mean Last Mean $p$-value Significant
Middle class 0.30-0.30 0.01-0.01 $7.06\times 10^{-73}$ Yes
Upper class 0.26-0.26 0.01-0.01 $1.25\times 10^{-58}$ Yes
Lower class 0.28-0.28 0.01-0.01 $4.33\times 10^{-66}$ Yes
Table 11: Changes in model confidence across social class groups, measured by the log probability of the selected Likert score between the first and final debate turns.
Political Ideology First Mean Last Mean $p$-value Significant
Libertarian-Right 0.29-0.29 0.01-0.01 $5.20\times 10^{-53}$ Yes
Libertarian-Left 0.30-0.30 0.01-0.01 $2.32\times 10^{-60}$ Yes
Authoritarian-Left 0.30-0.30 0.01-0.01 $1.26\times 10^{-44}$ Yes
Authoritarian-Right 0.22-0.22 0.01-0.01 $3.45\times 10^{-42}$ Yes
Table 12: Changes in model confidence across political ideology groups, measured by the log probability of the selected Likert score between the first and final debate turns.
Personality Trait First Mean Last Mean $p$-value Significant
High Extraversion 0.30-0.30 0.01-0.01 $2.38\times 10^{-24}$ Yes
High Conscientiousness 0.28-0.28 0.01-0.01 $4.88\times 10^{-22}$ Yes
Low Conscientiousness 0.23-0.23 0.02-0.02 $3.84\times 10^{-14}$ Yes
High Neuroticism 0.25-0.25 0.01-0.01 $3.34\times 10^{-20}$ Yes
High Openness 0.28-0.28 0.01-0.01 $2.10\times 10^{-16}$ Yes
Low Extraversion 0.29-0.29 0.01-0.01 $5.65\times 10^{-16}$ Yes
Low Agreeableness 0.35-0.35 0.00-0.00 $1.38\times 10^{-44}$ Yes
Low Neuroticism 0.26-0.26 0.02-0.02 $2.62\times 10^{-21}$ Yes
High Agreeableness 0.33-0.33 0.01-0.01 $2.90\times 10^{-18}$ Yes
Low Openness 0.19-0.19 0.01-0.01 $1.86\times 10^{-13}$ Yes
Table 13: Changes in model confidence across Big Five personality traits, measured by the log probability of the selected Likert score between the first and final debate turns.
Appendix O Debate Dynamics: Mode of Persuasion Score Changes
Table˜14 to Table˜19 present the changes in model’s mode of persuasions and their statistical significance, as measured by $t$-tests, across different persona dimensions.
Age Group Mode First Mean Last Mean p-value Significant
40 Ethos 2.65 2.54 0.02 Yes
40 Pathos 2.93 2.72 $1.3\times 10^{-4}$ Yes
40 Logos 3.48 3.35 $4.1\times 10^{-3}$ Yes
50 Ethos 2.64 2.43 $9.0\times 10^{-5}$ Yes
50 Pathos 3.07 2.86 $1.1\times 10^{-4}$ Yes
50 Logos 3.34 3.28 0.23 No
30 Ethos 2.53 2.41 $5.0\times 10^{-3}$ Yes
30 Pathos 3.03 2.80 $1.7\times 10^{-6}$ Yes
30 Logos 3.40 3.27 $3.9\times 10^{-4}$ Yes
20 Ethos 2.44 2.31 0.01 Yes
20 Pathos 2.95 2.75 $8.5\times 10^{-5}$ Yes
20 Logos 3.41 3.33 0.05 Yes
60 Ethos 2.68 2.49 $1.9\times 10^{-5}$ Yes
60 Pathos 3.03 2.68 $1.6\times 10^{-12}$ Yes
60 Logos 3.38 3.38 0.96 No
Table 14: Changes in model’s mode of persuasion across age groups between the first and final debate turns.
Age Group Mode First Mean Last Mean p-value Significant
Non-binary Ethos 2.54 2.39 $2.3\times 10^{-5}$ Yes
Non-binary Pathos 2.99 2.75 $4.6\times 10^{-10}$ Yes
Non-binary Logos 3.39 3.32 0.03 Yes
Male Ethos 2.60 2.43 $1.5\times 10^{-5}$ Yes
Male Pathos 2.94 2.70 $3.7\times 10^{-9}$ Yes
Male Logos 3.45 3.32 $1.6\times 10^{-5}$ Yes
Female Ethos 2.62 2.48 $2.6\times 10^{-4}$ Yes
Female Pathos 3.08 2.83 $3.0\times 10^{-10}$ Yes
Female Logos 3.36 3.33 0.36 No
Table 15: Changes in model’s mode of persuasion across gender groups between the first and final debate turns.
Age Group Mode First Mean Last Mean p-value Significant
India Ethos 2.47 2.32 $6.5\times 10^{-3}$ Yes
India Pathos 2.92 2.70 $7.3\times 10^{-5}$ Yes
India Logos 3.42 3.32 0.02 Yes
Brazil Ethos 2.68 2.55 $9.1\times 10^{-3}$ Yes
Brazil Pathos 3.09 2.81 $3.2\times 10^{-7}$ Yes
Brazil Logos 3.35 3.36 0.91 No
China Ethos 2.61 2.51 0.05 Yes
China Pathos 2.94 2.75 $7.1\times 10^{-4}$ Yes
China Logos 3.43 3.33 0.02 Yes
Nigeria Ethos 2.55 2.33 $3.6\times 10^{-5}$ Yes
Nigeria Pathos 2.94 2.69 $5.2\times 10^{-6}$ Yes
Nigeria Logos 3.39 3.33 0.19 No
France Ethos 2.62 2.46 $2.4\times 10^{-3}$ Yes
France Pathos 3.11 2.85 $1.0\times 10^{-5}$ Yes
France Logos 3.38 3.27 0.02 Yes
United States Ethos 2.61 2.46 $6.5\times 10^{-3}$ Yes
United States Pathos 3.04 2.77 $3.5\times 10^{-6}$ Yes
United States Logos 3.43 3.33 0.01 Yes
Table 16: Changes in model’s mode of persuasion across countries between the first and final debate turns.
Age Group Mode First Mean Last Mean p-value Significant
Middle class Ethos 2.58 2.40 $7.1\times 10^{-7}$ Yes
Middle class Pathos 2.98 2.74 $1.7\times 10^{-10}$ Yes
Middle class Logos 3.44 3.33 $2.1\times 10^{-4}$ Yes
Upper class Ethos 2.63 2.51 $1.5\times 10^{-3}$ Yes
Upper class Pathos 2.90 2.67 $5.0\times 10^{-9}$ Yes
Upper class Logos 3.45 3.39 0.03 Yes
Lower class Ethos 2.55 2.40 $5.6\times 10^{-5}$ Yes
Lower class Pathos 3.13 2.87 $4.8\times 10^{-10}$ Yes
Lower class Logos 3.30 3.26 0.13 No
Table 17: Changes in model’s mode of persuasion across social class groups between the first and final debate turns.
Age Group Mode First Mean Last Mean p-value Significant
Libertarian-Right Ethos 2.52 2.30 $9.0\times 10^{-8}$ Yes
Libertarian-Right Pathos 2.81 2.55 $1.8\times 10^{-9}$ Yes
Libertarian-Right Logos 3.52 3.41 $3.8\times 10^{-4}$ Yes
Libertarian-Left Ethos 2.58 2.38 $6.2\times 10^{-7}$ Yes
Libertarian-Left Pathos 3.22 3.03 $9.2\times 10^{-7}$ Yes
Libertarian-Left Logos 3.32 3.23 $5.9\times 10^{-3}$ Yes
Authoritarian-Left Ethos 2.61 2.50 0.03 Yes
Authoritarian-Left Pathos 3.13 2.87 $5.0\times 10^{-7}$ Yes
Authoritarian-Left Logos 3.34 3.26 0.03 Yes
Authoritarian-Right Ethos 2.64 2.57 0.12 No
Authoritarian-Right Pathos 2.84 2.56 $3.1\times 10^{-9}$ Yes
Authoritarian-Right Logos 3.42 3.41 0.86 No
Table 18: Changes in model’s mode of persuasion across political ideology groups between the first and final debate turns.
Age Group Mode First Mean Last Mean p-value Significant
High Extraversion Ethos 2.66 2.55 0.05 No
High Extraversion Pathos 2.99 2.90 0.17 No
High Extraversion Logos 3.42 3.39 0.66 No
High Conscientiousness Ethos 2.94 2.95 0.94 No
High Conscientiousness Pathos 2.78 2.56 $1.4\times 10^{-3}$ Yes
High Conscientiousness Logos 3.67 3.54 $5.9\times 10^{-3}$ Yes
Low Conscientiousness Ethos 2.33 2.10 $2.4\times 10^{-4}$ Yes
Low Conscientiousness Pathos 3.15 2.87 $8.2\times 10^{-6}$ Yes
Low Conscientiousness Logos 3.09 3.04 0.44 No
High Neuroticism Ethos 2.62 2.48 0.01 Yes
High Neuroticism Pathos 3.49 3.23 $1.4\times 10^{-5}$ Yes
High Neuroticism Logos 3.30 3.37 0.13 No
High Openness Ethos 2.73 2.55 $3.6\times 10^{-3}$ Yes
High Openness Pathos 3.14 2.99 0.06 No
High Openness Logos 3.38 3.41 0.61 No
Low Extraversion Ethos 2.68 2.57 0.09 No
Low Extraversion Pathos 3.10 2.78 $3.7\times 10^{-5}$ Yes
Low Extraversion Logos 3.39 3.32 0.24 No
Low Agreeableness Ethos 1.95 1.50 $1.3\times 10^{-18}$ Yes
Low Agreeableness Pathos 2.69 2.27 $1.2\times 10^{-13}$ Yes
Low Agreeableness Logos 3.43 3.22 $9.7\times 10^{-6}$ Yes
Low Neuroticism Ethos 2.84 2.78 0.21 No
Low Neuroticism Pathos 2.82 2.62 $3.8\times 10^{-3}$ Yes
Low Neuroticism Logos 3.57 3.52 0.29 No
High Agreeableness Ethos 2.93 2.96 0.59 No
High Agreeableness Pathos 3.43 3.40 0.65 No
High Agreeableness Logos 3.30 3.06 $8.3\times 10^{-5}$ Yes
Low Openness Ethos 2.72 2.66 0.34 No
Low Openness Pathos 2.77 2.43 $7.2\times 10^{-6}$ Yes
Low Openness Logos 3.40 3.40 0.94 No
Table 19: Changes in model’s mode of persuasion across Big Five personality traits between the first and final debate turns.
Appendix P Debate Dynamics: Debate Order
We randomly shuffled the speaking order of the two personas and conducted two-sided t-tests to evaluate whether speaking order had a statistically significant effect on debate confidence or consensus. The resulting p-values are reported in Table˜20. The tests fail to reject the null hypothesis that the distributions are the same under both speaking orders. In other words, debate order does not significantly affect model behavior in terms of confidence or consensus (particularly for GPT-4o), which appears highly robust to this variation.
Metric GPT-4o LLaMA-4-Maverick
First persona’s confidence score 0.89 0.17
Second persona’s confidence score 0.78 0.24
Consensus ratio 0.61 0.70
Table 20: P-values from two-sided t-tests assessing whether debate order (first vs. second speaker) significantly affects persona confidence scores and consensus ratio for GPT-4o and LLaMA-4-Maverick.